OpenStudy (anonymous):

Find the rational roots of x4 + 8x3 + 7x2 – 40x – 60 = 0 (1 point)2, 6 –6, –2 –2, 6 –6, 2

OpenStudy (anonymous):

please help me!!!!

OpenStudy (campbell_st):

you can use the factor theorem just evaluate f(2) f(6), f(-2) and f(-6) whenver you get a zero answer... the you have a rational root... e.g. f(2) = 2^4 + 8*2^3 + 7*2^2 - 40*2 - 60 = -32 so its not x = 2 try f(-2) = (-2)^4 + 8*(-2)^3 + 6*(-2)^2 - 40*(-2) - 60 = 0 so x = -2 is a rational zero try f(6) in the same manner... if it is zero then you have your 2 rational zeros... of not... then it must be -6 but just check hope this helps

OpenStudy (anonymous):

can u help me with another question?!

OpenStudy (campbell_st):

just post it in the normal fashoin and I'll look at it

OpenStudy (anonymous):

ok thank you